Matthew D. Fuller

Work History

2002– Independent Consultant, Madison, MS.

Responsibilities: Custom development of multiple software systems for data and business process management for multiple small-medium companies. Automation of existing procedures, co-discovery of new methods and capabilities. Ecommerce and payment tracking, interfacing with third-party API’s of various sorts. Principally web-based apps, implementing and managing the full stack. Specified, installed, and managed servers running that software. Managed servers providing additional web, mail, and other service hosting. Advised businesses on hardware and service purchases. Occasional additional misc IT work.

{2001–2002 \atop 1997–1999} Futuresouth Communications, Inc. / Graham Media Internet, Inc., Clarksdale, MS.

Responsibilities: Administered servers and networks for statewide ISP. Multiple services for access customers, web and mail hosting, colo provision. Datacenter management. Various hardware platforms for routing among multiple POPs and multiple upstreams, switching, remote access. Development of custom tools for technical and business management.

2000–2001 Linklabs Hosting, Inc., Cullman, AL.

Responsibilities: Implemented and managed web and mail hosting services. Adapted and built software packages to fulfill requirements. Handled technical-side management.

2000–2000 Linkfast Internet Services, Linkfast Inc., Cullman, AL.

Responsibilities: Administered servers and network for small local ISP. Usage monitoring and capacity planning. Telco interaction and provisioning. Network design and maintenance. Web, mail, and other service management. Assisted in logistics of transfer of customers and services to new owners upon sale of company.

CS/IT Experience

Programming

15+ years of diverse experience with software engineering and programming. Extensive back-end and full-stack web experience with PHP and Perl backended into SQL databases with HTML, CSS, and Javascript frontends. Significant Perl and C programming in the Unix environment. Scripting in support of system/network management using the above, Bourne shell, awk, etc. Some exposure to python and ruby. Passing knowledge of various others.

15+ years experience with relational database (SQL) utilization in support of existing and custom programs for data storage and analysis. Particular emphasis on PostgreSQL and MySQL, some knowledge of SQLite and others.

143 Shadow Hill Drive – Madison, MS 39110 601-853-3524 • fullermd@over-yonder.net https://www.over-yonder.net/^{∼} fullermd/ Extensive version control experience with Bazaar and CVS, with additional work- ing knowledge of git, Mercurial, SVN, monotone, and others.

Experience with build automation via make and cmake, as well as custom build systems. Experience with low-level debugging tools such as gdb and ktrace. Experience with documentation writing and markup using SGML, XML, TeX, POD, roff, markdown, and asciidoc. Some work with DSSSL and XSL.

Systems and Network Administration

15+ years experience with administration of Unix-like operating systems. Pri- mary emphasis on FreeBSD, moderate experience on various Linux distributions, and passing acquaintance with other Unices. Familiarity with all aspects, in- cluding system specification, design, and installation; hardware, network, and storage management; configuration management; ongoing maintainance upgrade management; backups and emergency preparation; transitions of live services; and emergency functions such as disaster recovery.

15+ years experience implementing and maintaining internal- and external- facing network services on Unix machines, such as HTTP, DNS, SMTP, POP, IMAP, NFS, SNMP, and others. Significant long-term experience with Apache, BIND, and Postfix. Experience with low-level debugging of wire protocols and server software.

7+ years experience specifying, designing, implementing, maintaining, and up- grading TCP/IP networks, both IPv4 and IPv6. Designing numbering and subnetting schemata, LAN and WAN management, VPN setups, telco/SP in- terfacing, and other mechanisms.

Various system and network monitoring tools like Graphite, MRTG, Cricket, spong, and others as well as custom tools. Specifying, installing, and maintain- ing dedicated network appliances, such as routers, switches, and remote access servers. Experience with Cisco, Hewlett-Packard, Lucent, Ascend, Livingston, and others.

Certifications

Past Cisco Systems CCNA (Cisco Certified Network Associate) and CCDA (Cisco Certified Design Associate) (now expired).

Independent Projects

Extensive long-term involvement in multiple open source projects, including FreeBSD and Bazaar, with more isolated contributions to many others. Main- tainer of ctwm window manager. Original open source code releases. Code contributions in various languages, bug reporting and triage, documentation, user support, community management, etc.

See my current Projects on my webpage .

143 Shadow Hill Drive – Madison, MS 39110 601-853-3524 • fullermd@over-yonder.net https://www.over-yonder.net/^{∼} fullermd/